ANALYSIS OF THE FIRST AND SECOND TRIADS OF H_{2}S FROM 2200 TO 4100 CM^{−1}

\underline{L. R. Brown}, J. A. Crisp, D. Crisp, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, California 91109; A. Perrin, Laboratoire de Physique Moléculaire Applications CNRS, Université Pierre et Marie Curie Bte 76, 4 Place Jussieu 75252 Paris Cedex 05, France; A. D. Bykov, O. V. Naumenko, M. A. Smirnov and L. N. Sinitsa, Institute of Atmospheric Optics, SB, Russian Academy of Science, 634055, Tomsk, Russia.

The 1800 - 5400 cm^{−1}region of H_{2}S was recorded at 0.011 cm^{−1}resolution using the McMath Fourier transform spectrometer located at Kitt Peak National Observatory. The positions of the first triad [2ν_{2}, ν_{1} and ν_{3}] near 4 μm and the second triad [3ν_{2}, ν_{1}+ν_{2} and ν_{2}+ν_{3}] near 2.7 μm were obtained for H_{2}^{32}S, H_{2}^{34}S and H_{2}^{33}S. The experimental upper states were fitted to the A-reduced Watson Hamiltonian to determine precise sets of rotational constants and Fermi and Coriolis coupling parameters. Line intensities of H_{2}^{32}S up to J = 17 and K_{a} = 12 were measured and modelled to ±2.5% using the 564 intensities of the first triad and 526 intensities of the second triad. The coefficients of the transformed moment expansion were obtained, corresponding to band strengths in cm^{−2}/atm at 296 K of: 0.33 for 2ν_{2}, 0.45 for ν_{1} and 0.12 for ν_{3}, 0.03 for 3ν_{2}, 1.82 for ν_{1}+ν_{3} and 2.87 for ν_{2}+ν_{3}.

Part of the research reported in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.